OpenStudy (anonymous):

Please Helppppppppp :D Solve 1/36 = 6x−3. x = −5 x = -3/2 x = 7/2 x = 1

OpenStudy (anonymous):

If a hospital patient is given 50 milligrams of medicine which leaves the bloodstream at 10% per hour, how many milligrams of medicine will remain in the system after 6 hours? Use the function A(t) = Iert. 12.39 mg 27.44 mg 45.24 mg 91.11 mg

HanAkoSolo (jamierox4ev3r):

kay... so if I understand correctly, its : \(\Large\frac{1}{36}=6x-3\)

HanAkoSolo (jamierox4ev3r):

for the first one

OpenStudy (anonymous):

Yes

HanAkoSolo (jamierox4ev3r):

okay..well, of you want to eliminate the fraction, what is a way to do that? Because that is how you are ultimately going to solve the equation

OpenStudy (anonymous):

Mm.. Multiply?

HanAkoSolo (jamierox4ev3r):

right, as that is the opposite of division. So what would you mutliply by?

OpenStudy (anonymous):

36?

OpenStudy (anonymous):

@raebaby420 :What is the question ? Do u need help ?

HanAkoSolo (jamierox4ev3r):

right, so then multiply everything in the equation by 36

HanAkoSolo (jamierox4ev3r):

so what would that give you?

OpenStudy (anonymous):

#1 @E.ali and sure.. more help the better

OpenStudy (anonymous):

Am I allowed to divide 1 by 36?

OpenStudy (anonymous):

Wait, what do you mean divide everything? How?

OpenStudy (anonymous):

Is it your question ?: 1/36=?????

OpenStudy (anonymous):

Whats the problem ???

OpenStudy (anonymous):

Up there is my question ^^^^ 1/36= 6x-3

HanAkoSolo (jamierox4ev3r):

wait what do you mean? all you do is this: then rewrite your equation based from that

OpenStudy (anonymous):

OK ! Just do this : 

OpenStudy (anonymous):

....huh?

OpenStudy (anonymous):

we have : 3+1/36=6x Just it :)

OpenStudy (anonymous):

-.- im so confused